OpenStudy (anonymous):

sin10+sin20+sin40+sin50=sin70+sin80.prove it..

OpenStudy (anonymous):

sin10+sin40+sin50+sin20=sin70+ sin80 Here we will use the formula sinA+sinB=2sin(A+B/2)cos(A-B/2) Take the left side of equation sin10+sin40+sin50+sin20 Arranging =(sin50+sin10)+(sin40+sin20) Applying the above formula. =2sin(50+10/2)cos(50-10/2)+2sin(40+20/… =2sin(30)cos(20)+2sin(30)cos(10) =2sin30{cos20+cos10} Again using the formula cosA+cosB= 2cos(A+B/2)cos(A-B/2) =2sin30{2cos(20+10/2).cos(20-10/2)} =2sin30{2cos(15).cos(5)} =2(1/2){2cos15.cos5} as sin30=1/2 =2cos15.cos5 Taking right side of equation sin70+ sin80 Using the formula sinA+sinB=2sin(A+B/2)cos(A-B/2) =2sin(70+80/2)cos(70-80/2) =2sin75cos5 =2sin(90-15)cos5 =2cos15.cos5 Hence proved